Jefferson Lab HALL A Logbook for August 99 - All EXCEPT Automatic Entries

  • Main INDEX

    Tuesday

  • 26688 08/31/99 23:10 Roche status of scintillator
  • 26675 08/31/99 16:19 Eugene Moller result
  • 26654 08/31/99 07:55 acochran Shift Summary
  • 26651 08/31/99 05:52 acochran TS1 not responding
  • 26645 08/31/99 01:49 acochran DAQ glitch
  • Monday

  • 26639 08/30/99 23:59 R Roche Shift summary
  • 26633 08/30/99 22:02 R. Michaels run 1485 partially lost
  • 26632 08/30/99 21:53 R Roche 8.1.18 HV trip
  • 26616 08/30/99 15:33 perdrisat ROC14 dead
  • 26613 08/30/99 13:02 perdrisat back to LD2
  • 26612 08/30/99 12:51 perdrisat raster cannot be turned off
  • 26611 08/30/99 12:47 perdrisat helicity -current correlation
  • 26610 08/30/99 12:43 perdrisat Moeller results
  • 26605 08/30/99 12:03 Eugene Moller results
  • 26604 08/30/99 09:45 perdrisat FPP X2 has a hole
  • 26603 08/30/99 09:26 perdrisat FPP HV 8.1.18
  • 26599 08/30/99 09:11 perdrisat FPP HV 8.1.18
  • 26592 08/30/99 07:05 markowitz gas trip off followup
  • 26589 08/30/99 06:10 markowitz gas system trip
  • 26577 08/30/99 00:51 tree scintillator peak position
  • Sunday

  • 26572 08/29/99 23:26 Adaq TDC of SPEC_H.S2_L3 has some grass
  • 26569 08/29/99 22:17 X. jiang, S. Looked at run 1603-1605
  • 26538 08/29/99 07:16 tree FPP 4 tripped off
  • 26533 08/29/99 05:59 tree Hadron arm angle?
  • 26532 08/29/99 05:04 tree FPP chamber 4 trip
  • 26522 08/29/99 00:35 acochran Junk Runs 1538 to 1541
  • Saturday

  • 26521 08/28/99 21:53 Calarco CHL crash
  • 26520 08/28/99 21:26 Calarco CHL crash
  • 26517 08/28/99 21:05 Calarco HV
  • 26516 08/28/99 20:58 Calarco FPP HV trip
  • 26513 08/28/99 20:31 Calarco HV
  • 26506 08/28/99 17:43 R. Michaels I/O errors on /adaqs2/data4
  • 26498 08/28/99 14:56 Perdrisat ER1 error
  • 26495 08/28/99 14:25 Perdrisat FPP HV
  • 26493 08/28/99 13:49 Perdrisat Moller quads
  • 26484 08/28/99 12:07 GARIBALDI Probl;ems with DAQ
  • 26483 08/28/99 11:50 R. Michaels "corrupt" event
  • 26481 08/28/99 11:23 R. Michaels Re: ROC14 problems
  • 26480 08/28/99 11:20 R. Michaels bug in my data handling fixed
  • 26474 08/28/99 10:34 M Jones trigger downloaded
  • 26472 08/28/99 10:21 R. Michaels root partition - big deadtime
  • 26470 08/28/99 09:54 M JOnes coda problem
  • 26469 08/28/99 09:53 M Jones roc 14 hung
  • 26462 08/28/99 08:52 GARIBALDI Y-TARGET SIEVE
  • 26403 08/28/99 01:57 markowitz target change
  • Friday

  • 26362 08/27/99 20:37 Eugene Moller results
  • 26361 08/27/99 16:41 mitchell Injector work
  • 26355 08/27/99 15:53 Woo target change
  • 26339 08/27/99 13:16 Woo Target Change
  • 26336 08/27/99 12:14 Woo target change
  • 26323 08/27/99 10:42 M Jones hadron camera
  • 26309 08/27/99 08:51 Woo Target Change
  • 26276 08/27/99 03:09 Keppel target fan speed
  • 26271 08/27/99 01:13 markowitz changed trigger
  • 26263 08/27/99 00:00 R Roche Shift summary
  • Thursday

  • 26261 08/26/99 23:30 R Roche Run #1472 Junk
  • 26248 08/26/99 18:22 R Roche Magnet summary page
  • 26245 08/26/99 17:41 Eugene Moller results
  • 26244 08/26/99 16:19 M Jones SHift summary
  • 26239 08/26/99 10:36 Adaq Power Glitch
  • 26238 08/26/99 08:40 Woo power glitch
  • 26236 08/26/99 08:18 Meekins BIG TIME CHL CRASH
  • 26235 08/26/99 08:08 Meekins Shift Summary:
  • 26228 08/26/99 05:08 Meekins FPP ancd VDC gas flow
  • 26223 08/26/99 03:38 Meekins Bremsstrahlung Radiator Control
  • 26220 08/26/99 02:30 Meekins Moved Beam spot
  • 26214 08/26/99 00:15 Meekins FPP and VDC gas flow
  • Wednesday

  • 26201 08/25/99 18:07 M Jones See some noise in the VDCs
  • 26199 08/25/99 17:27 M Jones 8.1.18 reset
  • 26190 08/25/99 15:47 Stevensm BCM meter calibration
  • 26189 08/25/99 15:37 Folts Q1 fix
  • 26182 08/25/99 12:37 M Jones After recesiation current 45.5uA
  • 26176 08/25/99 11:44 M Jones 8.1.18 tripped
  • 26174 08/25/99 11:07 M Jones Status of source
  • 26164 08/25/99 08:01 Meekins Shift summary
  • 26150 08/25/99 02:40 Meekins electron arm Q1 trip
  • 26147 08/25/99 02:36 Meekins DAQ crash
  • 26144 08/25/99 02:19 Meekins Shift Summary
  • 26135 08/25/99 00:29 X. jiang spectronic screen at run 1422.
  • Tuesday

  • 26117 08/24/99 22:49 J. Hovdebo Spec settings: 60 degree cm 25 MeV offset
  • 26095 08/24/99 19:15 Eugene Moller results
  • 26094 08/24/99 18:29 Nilanga 1.66 Beam energy is back
  • 26091 08/24/99 17:12 Krishni day shift summary
  • 26090 08/24/99 17:11 M Jones trigger setup file downloaded
  • 26088 08/24/99 15:34 X. jiang junk picture.
  • 26087 08/24/99 11:45 X. Jiang hallA went into beam permit.
  • 26086 08/24/99 11:42 X. Jiang FPP High Voltage Arrangement
  • 26084 08/24/99 09:28 M Jones chamber gas flow
  • 26083 08/24/99 08:55 Eugene Moller results
  • 26081 08/24/99 06:29 Meekins FPP and VDC gas flow
  • 26080 08/24/99 06:06 Meekins Hall C access
  • 26078 08/24/99 05:39 Meekins power glitch
  • 26076 08/24/99 04:24 Meekins VDC anf FPP gas flow
  • 26075 08/24/99 04:16 Meekins VDC multiplicities for run 1400
  • 26067 08/24/99 02:30 Meekins FPP time spectra for run 1398
  • 26066 08/24/99 02:28 Meekins VDC multiplicities for run 1398
  • 26064 08/24/99 02:05 Meekins FPP and VDC gas flow settings
  • 26059 08/24/99 00:08 Meekins Gas flow for FPP and VDC
  • Monday

  • 26056 08/23/99 23:18 J. Hovdebo Spec Settings: 75 degree cm 35 MeV offsets
  • 26051 08/23/99 21:21 Nilanga FPP HV reduced futher
  • 26044 08/23/99 18:12 Nilanga FPP HV reduced
  • 26039 08/23/99 15:59 glashausser day journal continued
  • 26036 08/23/99 15:25 R. Michaels changing kine to 75 CM, 35 MeV
  • 26034 08/23/99 14:39 R. Michaels network and hac ok now
  • 26033 08/23/99 14:08 R. Michaels network problems
  • 26028 08/23/99 12:44 mitchell Computer problems, Network
  • 26023 08/23/99 10:42 R. Michaels espacelogon
  • 26022 08/23/99 10:36 R. Michaels new dhist ref histograms
  • 26019 08/23/99 09:37 X. Jiang Instructions of Online Replay
  • 26018 08/23/99 09:33 R. Michaels FPP trip
  • 26015 08/23/99 09:06 glashausser FPP trip
  • 26007 08/23/99 06:15 Meekins Gas flow settings
  • 26004 08/23/99 05:42 Meekins Gas flow settings
  • 26003 08/23/99 05:01 Meekins FPP TDC spectra for rear chambers run 1376
  • 26002 08/23/99 05:00 Meekins FPP tdc spectra for run 1376 front chambers
  • 26001 08/23/99 04:52 Meekins VDC multiplicities for run 1375
  • 26000 08/23/99 04:37 Meekins FPP and VDC gas flow settings
  • 25997 08/23/99 04:25 Meekins FPP straw eff for rear chambers for run 1375
  • 25996 08/23/99 04:24 Meekins FPP straw efficiencies for run 1375 for front chambers
  • 25995 08/23/99 04:14 Meekins FPP tdc spectra for run 1375 rear chambers
  • 25994 08/23/99 04:13 Meekins FPP TDC spectra for run 1375/ Chnaged FPP HV on 7.0.0 and 7.0.1
  • 25993 08/23/99 03:59 Meekins FPP and VDC gas flow settings
  • 25992 08/23/99 03:39 Meekins Rear FPP TDC spectra
  • 25990 08/23/99 03:35 Meekins FPP TDC spectra and Voltage Settings
  • 25987 08/23/99 03:00 Meekins FPP and VDC gas flow settings
  • 25986 08/23/99 02:59 Meekins Run 1373 bad FPP data
  • 25985 08/23/99 02:43 Meekins FPP high voltage trip
  • 25984 08/23/99 02:34 Meekins FPP and VDC gas system
  • 25983 08/23/99 02:14 Kumbartzki FPP HV 8.1.18 off
  • 25982 08/23/99 02:11 Meekins FPP high voltage left off
  • 25981 08/23/99 02:09 Meekins ROC reboot
  • 25978 08/23/99 01:44 meekins Inconsistency with Hall A gas system ethane bottles
  • Sunday

  • 25969 08/22/99 23:55 glashausser swing shift journal
  • 25966 08/22/99 23:01 glashausser mag settings at end of 105
  • 25956 08/22/99 20:55 mitchell Spontaneous IOC reboot
  • 25953 08/22/99 20:15 Dumalski Multiplicity run #1362
  • 25952 08/22/99 20:12 Dumalski FPP drift times
  • 25951 08/22/99 20:11 Dumalski VDC multiplivity run # 1362
  • 25950 08/22/99 20:08 glashausser VDC efficiency
  • 25949 08/22/99 20:08 Dumalski VDC Efficiency run #1362
  • 25946 08/22/99 19:29 Dumalski fpp straw efficiency
  • 25945 08/22/99 19:26 Dumalski fpp straw efficiency
  • 25942 08/22/99 18:47 Dumalski VDC efficiency run 1359
  • 25941 08/22/99 18:32 mitchell Target IOC Reboot
  • 25940 08/22/99 18:26 Dumalski VDC Multiplicity Check
  • 25937 08/22/99 17:44 glashausser FPP currents OK
  • 25935 08/22/99 17:01 glashausser TV monitor on e arm
  • 25934 08/22/99 16:21 glashausser FPP trip
  • 25932 08/22/99 15:53 eschulte Day Shift Log
  • 25929 08/22/99 15:44 Z. Chai FPP Straw efficiency ( Rear Chamber)
  • 25928 08/22/99 15:43 Z. Chai FPP Straw Efficiency (Front Chamber)
  • 25927 08/22/99 15:24 eschulte 15:25 FPP current check
  • 25924 08/22/99 14:54 Z. Chai VDC Wire efficiency Run #1354
  • 25920 08/22/99 14:35 Z. Chai VDC multiplicity Run# 1255
  • 25919 08/22/99 14:29 Z. Chai VDC efficiency Check
  • 25918 08/22/99 14:25 Z. Chai VDC multiplicity Run# 1354
  • 25917 08/22/99 14:21 eschulte 14:20 FPP check
  • 25916 08/22/99 13:59 Z. Chai FPP Straw Efficiency ( Rear Chamber)
  • 25915 08/22/99 13:58 Z. Chai FPP Straw efficiency ( Front Chamber)
  • 25912 08/22/99 13:33 eschulte 13:30 FPP current check
  • 25909 08/22/99 12:35 eschulte 12:30 FPP check
  • 25906 08/22/99 11:32 Z. Chai FPP Straw efficiency ( Rear Chamber)
  • 25905 08/22/99 11:31 Z. Chai FPP Straw efficiency (Front Chamber)
  • 25904 08/22/99 11:21 Z. Chai VDC Multiplicity Run#1351
  • 25901 08/22/99 11:09 eschulte 11:10 FPP check
  • 25899 08/22/99 10:21 eschulte 10:20 FPP current check
  • 25898 08/22/99 09:49 Z. Chai VDC Efficiency
  • 25896 08/22/99 09:47 eschulte 105 cm 70 MeV offset magnet screen
  • 25895 08/22/99 09:46 eschulte Electron arm camera
  • 25894 08/22/99 09:44 eschulte 09:00 FPP current check
  • 25893 08/22/99 09:39 eschulte slight argon adjustment
  • 25892 08/22/99 09:26 Z. Chai FPP straw efficiency
  • 25891 08/22/99 09:25 Z. Chai FPP straw efficiency
  • 25887 08/22/99 08:34 L.Bimbot Cryotarget
  • 25886 08/22/99 08:15 L.Bimbot Cryotarget
  • 25881 08/22/99 07:50 R. Roche Run #1347 Multiplicities check
  • 25880 08/22/99 07:49 R. Roche Run #1347 FPP check
  • 25879 08/22/99 07:48 R. Roche Run #1347 VDC check
  • 25872 08/22/99 06:12 R. Roche Run #1344 Muliplicities check
  • 25871 08/22/99 06:11 R. Roche Run #1344 FPP check
  • 25870 08/22/99 06:10 R. Roche Run #1344 VDC check
  • 25865 08/22/99 05:18 R. Roche Target IOC reboot
  • 25864 08/22/99 05:08 R. Roche Run #1340 FPP rear chamber efficiency
  • 25863 08/22/99 05:07 R. Roche Run #1340 FPP front chamber efficiency
  • 25860 08/22/99 04:41 R. Roche Run #1340 Multiplicities Check
  • 25859 08/22/99 04:40 R. Roche Run #1340 FPP check
  • 25857 08/22/99 04:39 R. Roche Run #1340 VDC check
  • 25844 08/22/99 02:08 R. Roche Run #1334 Multiplicities check
  • 25843 08/22/99 02:07 R. Roche Run #1334 FPP check
  • 25842 08/22/99 02:06 R. Roche Run #1334 VDC check
  • 25835 08/22/99 01:07 R. Roche Run #1329 FPP rear chamber straw efficiency
  • 25834 08/22/99 01:05 R. Roche Run #1329 FPP straw-by-straw efficiency front chamber
  • 25829 08/22/99 00:29 R. Roche Multiplities Check
  • 25828 08/22/99 00:28 R. Roche FPP check
  • 25827 08/22/99 00:26 R. Roche VDC check
  • Saturday

  • 25823 08/21/99 23:58 glashausser swing shift journal
  • 25808 08/21/99 21:59 Dumalski run 1318 check Multiplities
  • 25807 08/21/99 21:58 Dumalski run 1318 fpp check -- drift times
  • 25806 08/21/99 21:56 Dumalski run 1318 VDC check
  • 25800 08/21/99 21:17 J. Hovdebo Spec settings 120 degree cm 70 MeV offset
  • 25790 08/21/99 19:02 Dumalski run 1313 fpp check -- drift times
  • 25789 08/21/99 18:47 Dumalski run 1313 check Multiplities
  • 25788 08/21/99 18:43 glashausser fpp trip
  • 25785 08/21/99 18:37 Dumalski run 1313 VDC check
  • 25783 08/21/99 18:03 J. Hovdebo ep false assymetry for 1.667 GeV #3
  • 25774 08/21/99 15:59 eschulte Day Shift Log
  • 25773 08/21/99 15:40 eschulte run 1305 detector check: FPP (cont.)
  • 25772 08/21/99 15:39 eschulte run 1305 efficiencies: FPP
  • 25771 08/21/99 15:38 eschulte run 1305 efficiencies: VDCs
  • 25768 08/21/99 15:26 eschulte 15:30 FPP current check
  • 25765 08/21/99 14:29 eschulte run 1305 fpp check--multiplicities
  • 25764 08/21/99 14:28 Adaq run 1305 fpp check--drift times
  • 25759 08/21/99 14:25 Adaq run 1305 VDC check
  • 25760 08/21/99 14:25 eschulte 14:20 FPP current check
  • 25756 08/21/99 13:50 eschulte setup 2 magnet screen
  • 25755 08/21/99 13:41 M Jones vdc eff
  • 25754 08/21/99 13:38 eschulte 13:30 currents on FPP
  • 25751 08/21/99 13:13 M Jones fpp rear chamber straw eff
  • 25750 08/21/99 13:10 M Jones FPP straw by straw efficiency front chamber
  • 25746 08/21/99 11:56 eschulte run 1302 fpp check--multiplicities
  • 25745 08/21/99 11:55 eschulte run 1302 fpp check--drift times
  • 25744 08/21/99 11:53 eschulte run 1302 VDC check
  • 25741 08/21/99 11:01 eschulte a picture of tc-cor for today
  • 25740 08/21/99 10:58 eschulte run 1301 fpp check
  • 25739 08/21/99 10:55 eschulte run 1301 fpp drift times
  • 25738 08/21/99 10:52 eschulte 1301 VDC's
  • 25737 08/21/99 10:36 X. Jiang Fpp drift time are fine in run 1298.
  • 25736 08/21/99 10:33 X. Jiang FPP are working fine.
  • 25735 08/21/99 10:12 X. Jiang VDCs are almost recovered.
  • 25723 08/21/99 07:59 S. Strauch shift summary
  • 25721 08/21/99 07:24 krishniw vdc multiplicity.
  • 25719 08/21/99 06:53 krishniw vdc multiplicity
  • 25716 08/21/99 05:30 Krishniw vdc multiplicity
  • 25708 08/21/99 03:19 Meekins Changed the Gas flow for VDCs
  • 25707 08/21/99 03:00 S. Strauch magnets settings ep for 1.667 GeV d(gp), false asym - 1
  • Friday

  • 25704 08/20/99 23:56 glashausser swing shift summary
  • 25703 08/20/99 23:39 X. Jiang test with VDC threshold voltage set to 2.5 volts. No hope there.
  • 25702 08/20/99 23:35 X. Jiang slight improvement in VDC.
  • 25701 08/20/99 23:09 Meekins Change of VDC gas system flows
  • 25694 08/20/99 21:20 punjabi checklist pg3
  • 25693 08/20/99 21:11 punjabi checklist pg1,2
  • 25692 08/20/99 19:20 Meekins VDC gas system flow settings
  • 25691 08/20/99 19:18 Meekins Changes to the VDC gas sytem
  • 25682 08/20/99 13:13 eschulte 1283 VDC
  • 25680 08/20/99 12:21 eschulte 1283 VDC's
  • 25677 08/20/99 11:44 Nilanga Beta optimization improved
  • 25676 08/20/99 11:29 eschulte 1281 VDC's
  • 25675 08/20/99 11:18 Adaq Beta problems solved
  • 25667 08/20/99 08:51 eschulte 08:50 am power glitch
  • 25664 08/20/99 08:20 M Jones owl shift history
  • 25647 08/20/99 03:27 Louis BIMBOT magnets 08/20/99 @3.30 a.m.
  • 25636 08/20/99 00:31 Eugene Moller results
  • 25635 08/20/99 00:12 A.J. Sarty shift summary
  • Thursday

  • 25633 08/19/99 21:54 sarty VDC/FPP gas mixing
  • 25630 08/19/99 17:22 X. jiang check VDC E-arm.
  • 25629 08/19/99 17:18 X. jiang H-arm VDC.
  • 25624 08/19/99 16:01 J.P. Chen Day shift summary
  • 25621 08/19/99 12:53 Brash New database and online kumac
  • 25620 08/19/99 12:07 Saha Arc-energy Correction
  • 25613 08/19/99 11:14 Saha Arc-Energy measurement
  • 25611 08/19/99 09:53 X. Jiang Gas problem again this morning.
  • 25609 08/19/99 08:20 eschulte 60 degree cm gamma p magnets
  • 25608 08/19/99 08:01 S. Strauch Shift summary
  • 25605 08/19/99 07:02 S. Strauch spectrometer screen, run #1256
  • 25603 08/19/99 06:31 S. Strauch spectrometer screen, run #1255
  • 25590 08/19/99 01:42 S. Strauch gamma-p, 60 deg. settings
  • Wednesday

  • 25573 08/18/99 23:45 L.Bimbot swing shift history
  • 25569 08/18/99 23:36 R Roche Run 1240 FPP efficiency
  • 25570 08/18/99 23:36 sarty new magnet settings (starting run 1238)
  • 25561 08/18/99 22:24 X. jiang spectrometer screen at run 1238
  • 25560 08/18/99 22:19 X. jiang carbon door setting at 1238.
  • 25557 08/18/99 21:06 sarty wrong trigger/ strange rates / Q1(elec) trip!
  • 25552 08/18/99 20:41 R.Roche Hot wire in FPP U2
  • 25550 08/18/99 20:32 sarty raster check
  • 25549 08/18/99 20:27 sarty magnet settings - ep false asym check
  • 25541 08/18/99 19:30 saha harps checked & working
  • 25533 08/18/99 15:25 segal Gas flow problems.
  • 25532 08/18/99 14:17 R. Michaels self timing looks good too
  • 25531 08/18/99 14:12 R. Michaels coincidence time ok
  • 25530 08/18/99 13:57 Eugene Moller results
  • 25526 08/18/99 12:24 X. Jiang Moller Q2 and Q3 has been off.
  • 25525 08/18/99 11:50 eschulte Magnet screen for ep/GEp check
  • 25518 08/18/99 10:59 mitchell Daily Power Glitch
  • 25515 08/18/99 10:35 mitchell Beam Spot
  • 25498 08/18/99 08:03 S. Strauch setup change (90 deg, 40 MeV, off)
  • 25494 08/18/99 07:58 S. Strauch Shift summary
  • 25493 08/18/99 07:51 S. Strauch E-arm, Q1 trip
  • 25491 08/18/99 07:21 Adaq double peak in coincidence time
  • 25488 08/18/99 06:56 Adaq double peak in coincidence time since run 1209
  • 25487 08/18/99 06:42 S. Strauch setup change (ep for 1.15 GeV, #3)
  • 25478 08/18/99 05:02 S. Strauch setup change (ep for 1.15 GeV, #2)
  • 25473 08/18/99 03:57 S. Strauch setup change (ep for 1.15 GeV, #1)
  • 25462 08/18/99 01:52 S. Strauch E-Arm, Q1 trip
  • 25459 08/18/99 01:40 S. Strauch setup change
  • Tuesday

  • 25448 08/17/99 23:57 X. Jiang 8/17/98 swing shift summary
  • 25441 08/17/99 22:40 C. C. Chang Spectrometer settings
  • 25435 08/17/99 21:46 C. C. Chang GEp 4th kinematics
  • 25428 08/17/99 20:44 C. C. Chang run # 1184 third elast. ep
  • 25421 08/17/99 19:20 C. C. Chang pe=1.277, theta_e=34.09,ph=0.961, theta_h=49.591
  • 25416 08/17/99 18:15 x. Jiang em vs pm in run 1178
  • 25413 08/17/99 17:44 C. C. Chang Magnets screen dump
  • 25406 08/17/99 16:54 X. Jiang H-arm s2 adc
  • 25404 08/17/99 16:49 C. C. Chang Magnets screen
  • 25400 08/17/99 16:13 X. Jiang changed HV of H-S2-R2
  • 25397 08/17/99 16:02 eschulte Day Shift Log
  • 25387 08/17/99 15:05 R. Michaels ped suppression off for helicity data
  • 25385 08/17/99 14:53 eschulte 120 degree CM gammap magnet page
  • 25380 08/17/99 13:04 eschulte New door position
  • 25375 08/17/99 10:59 R. Michaels mods to DAQ scripts
  • 25367 08/17/99 09:17 eschulte new magnet screen
  • 25360 08/17/99 07:53 bimbot owl shift 8/17 : history
  • 25344 08/17/99 01:41 sarty changed raster (again) to square (run 1149)
  • 25339 08/17/99 00:33 sarty new raster size (run 1147)
  • 25335 08/17/99 00:18 sarty incorrect raster size (run 1146)
  • 25333 08/17/99 00:01 jgao swing shift summary
  • Monday

  • 25324 08/16/99 20:50 Adaq spectrometer setting for e-arm 14.938 degree, P_0=2.104 GeV/c
  • 25314 08/16/99 15:55 mitchell shift summary
  • 25309 08/16/99 15:29 mitchell ESR flow oscillations
  • 25306 08/16/99 14:53 mitchell Fast Feedback-On
  • 25305 08/16/99 14:12 eschulte Shift checklist 8/13-8/15 Page 3
  • 25304 08/16/99 14:09 eschulte Shift checklists 8/13 through 8/15 Pages 1&2
  • 25303 08/16/99 14:06 mitchell Beam back-RF problems
  • 25301 08/16/99 12:02 mitchell Beam Down for RF problems
  • 25299 08/16/99 11:16 mitchell 15 cm D2 target
  • 25294 08/16/99 10:08 mitchell Fast feedback test are over
  • 25293 08/16/99 10:07 mitchell Target operator incompetence
  • 25291 08/16/99 09:39 mitchell Invasive Fast Feedback
  • 25289 08/16/99 08:52 mitchell Fast Feedback - Test
  • 25287 08/16/99 08:35 mitchell Switch to D2 4cm target
  • 25283 08/16/99 08:07 mitchell q1 electron arm down again
  • 25281 08/16/99 07:35 Nilanga Radiator out
  • 25276 08/16/99 06:41 Nilanga Run 1126
  • 25275 08/16/99 06:24 Nilanga Halog Entries for rusn 1119-1125
  • 25273 08/16/99 06:17 Nilanga LD2 4cm target runs.
  • 25271 08/16/99 05:17 Nilanga Run 1183
  • 25268 08/16/99 05:15 Nilanga Run 1184
  • 25266 08/16/99 05:08 Nilanga Cryo IOC crash
  • 25262 08/16/99 04:13 Crovelli electron magnets cryogen problem
  • 25261 08/16/99 04:01 ddutta Electron arm S1 ADCs
  • 25257 08/16/99 01:41 Crovelli Beam down
  • Sunday

  • 25250 08/15/99 23:58 jgao swing shift summary
  • 25247 08/15/99 23:37 Bimbot Cryotarget
  • 25242 08/15/99 21:12 Bimbot Cryotarget
  • 25241 08/15/99 20:48 jgao spectrometer settings for e-arm 143 degree cm, h-arm 90 degree cm
  • 25216 08/15/99 16:05 Z. Chai Target Problem
  • 25215 08/15/99 16:00 eschulte Day Shift Log
  • 25210 08/15/99 13:57 eschulte Strange readings on gas flows
  • 25186 08/15/99 08:02 M Jones Shift summary
  • 25181 08/15/99 07:02 ddutta IOC reboot
  • 25173 08/15/99 05:01 Crovelli configchange2
  • 25161 08/15/99 01:40 Crovelli Configchange
  • 25155 08/15/99 00:54 M Jones 8.1.18 tripped off decided to leave off
  • 25154 08/15/99 00:18 M Jones Reset 8.1.18 before 1075
  • Saturday

  • 25150 08/14/99 23:57 jgao swing shift summary
  • 25146 08/14/99 23:16 sarty remove radiator
  • 25145 08/14/99 23:15 sarty cryotarget IOC reboot
  • 25141 08/14/99 22:14 sarty radiator move in (6%)
  • 25139 08/14/99 22:07 dumalski e arm 53 cm, h arm 90 cm
  • 25138 08/14/99 22:02 sarty remove radiator
  • 25132 08/14/99 21:15 jgao Spectrometer setting H arm: 90 degree cm 70 MeV offset E arm: 37 degree cm
  • 25131 08/14/99 21:15 sarty cryotarget move to hydrogen
  • 25126 08/14/99 20:06 sarty radiator move
  • 25122 08/14/99 19:54 jgao spectrometer setting for E-arm : 37 degree cm, H-arm: 90 degree cm, 40 MeV offset
  • 25118 08/14/99 19:19 x. jiang carbon door setting at run 1053
  • 25117 08/14/99 19:03 Eugene Moller results
  • 25114 08/14/99 16:20 Jordan e arm 37 degree cm h arm 90 degree cm 70 meV offset
  • 25111 08/14/99 16:09 Holt day shift
  • 25109 08/14/99 16:02 bimbot cryotarget
  • 25106 08/14/99 14:29 Bimbot Cryotarget to sleep
  • 25105 08/14/99 12:32 Bimbot cryotarget
  • 25104 08/14/99 12:22 Bimbot Cryotarget
  • 25094 08/14/99 08:06 Kelly shift summary
  • 25086 08/14/99 05:39 Kelly brief beam outage
  • 25081 08/14/99 04:22 Kelly frequent trips on 8.1.18
  • 25076 08/14/99 03:04 Kelly trips
  • 25073 08/14/99 01:45 Kelly FPP trip
  • 25072 08/14/99 01:42 Kelly deuterium target problem
  • 25067 08/14/99 00:53 x. jiang checked fpp door,
  • 25062 08/14/99 00:38 Kelly 6 in door
  • 25055 08/14/99 00:04 Adaq Spectrometer Settings 90 degree cm 40MeV offset
  • 25054 08/14/99 00:02 jpchen/jgao Swing shift summary
  • Friday

  • 25050 08/13/99 23:38 perdrisat FPP HV reset
  • 25041 08/13/99 22:23 perdrisat FPP HV reset
  • 25037 08/13/99 21:47 perdrisat FPP HV reset
  • 25023 08/13/99 20:08 Adaq spectrometer setting 45 degree cm 25 MeV
  • 25018 08/13/99 19:15 Adaq spectrometer setting for 60 degree cm, 25 MeV (kinematics #7)
  • 25011 08/13/99 18:35 Holt day shift summary (15-16)
  • 25004 08/13/99 17:16 jpchen Moller Q2 power supplies
  • 24983 08/13/99 15:45 Holt Target posts block scattered beam
  • 24982 08/13/99 15:45 eschulte Shift checklists--8-12 through 8-13 p. 3
  • 24981 08/13/99 15:43 eschulte Shift Checklists--8-12 through 8-13 pp. 1 and 2
  • 24974 08/13/99 15:10 gilman h+ gated on h+
  • 24973 08/13/99 15:04 X. jiang, R. ix the FPP ch4 hv card-u5 problem.
  • 24972 08/13/99 13:45 Holt Day shift, partial summary
  • 24968 08/13/99 12:42 M Jones checked runend.tcl
  • 24936 08/13/99 09:28 eschulte Data recording problems
  • 24925 08/13/99 07:58 Crovelli Configchange
  • 24923 08/13/99 07:52 Kelly shift summary
  • 24913 08/13/99 05:12 Meekins FPP HV trip
  • 24906 08/13/99 04:23 Crovelli tripped7120
  • 24902 08/13/99 04:02 Kelly incorrect door for dummy run
  • 24899 08/13/99 03:47 Crovelli Configuration 3:40am
  • 24888 08/13/99 02:10 Kelly high dead time at end of 967
  • 24885 08/13/99 01:42 Kelly beam restored
  • 24880 08/13/99 00:28 meekins Beam down for hall C energy change
  • Thursday

  • 24872 08/12/99 23:37 mitchell Problem in End of Run Entries
  • 24858 08/12/99 21:56 kuss HowTo set the trigger file
  • 24820 08/12/99 17:43 x jiang spectrometer screen at the end of run 931.
  • 24806 08/12/99 16:33 X. jiang moved in the 4th foil of the raditor
  • 24803 08/12/99 16:26 M JOnes New carbon door GUI
  • 24797 08/12/99 16:07 Ransome day journal
  • 24786 08/12/99 14:57 Eugene Moller results
  • 24785 08/12/99 13:48 Ransome fpp carbon doors - confusion
  • 24784 08/12/99 11:15 M Jones added targ position in runstart
  • 24770 08/12/99 07:55 Kelly shift summary
  • 24763 08/12/99 07:26 simon magnet gui for +2%
  • 24758 08/12/99 06:42 simon magnet gui for -2%
  • 24750 08/12/99 05:41 simon magnet gui for -4%
  • 24749 08/12/99 05:37 Kelly no gas flow to VDCs
  • 24742 08/12/99 04:50 Kelly adjust p0
  • 24740 08/12/99 04:30 Adaq coda cold start
  • 24737 08/12/99 04:23 simon daq quitting/restarting
  • Wednesday

  • 24722 08/11/99 09:10 kuss OTR monitor usage directions
  • 24721 08/11/99 07:28 Kelly shift summary
  • Tuesday

  • 24719 08/10/99 23:50 Crovelli webjunk
  • 24718 08/10/99 23:49 Crovelli webjunknick
  • 24716 08/10/99 23:38 Crovelli webstuffjunk
  • 24715 08/10/99 19:26 X. jiang RF timing signal
  • 24712 08/10/99 13:44 eschulte Gas Panel Alarm
  • 24709 08/10/99 10:32 gilman RFtime,FPP HV, helicity scaler
  • 24708 08/10/99 08:06 eschulte Power glitch and target problems
  • 24707 08/10/99 07:47 Kooijman HV tripped
  • Monday

  • 24706 08/09/99 23:55 bimbot Cryotarget
  • 24705 08/09/99 20:01 x. jiang changes on FPP chamber4 HV cabeling.
  • 24702 08/09/99 18:51 M Jones Carbon door GUI
  • 24697 08/09/99 06:55 Morand Shift summary
  • 24685 08/09/99 01:33 x. jiang run 883, 6% radiator in.
  • 24682 08/09/99 01:14 ransome run plan
  • 24681 08/09/99 01:04 x. jiang radiator out run
  • 24678 08/09/99 00:49 Kooijman check coincidence timing
  • 24677 08/09/99 00:43 Bimbot cryotarget lifting
  • 24676 08/09/99 00:34 Ransome shift plan for night
  • 24674 08/09/99 00:31 Ransome Hadron motion restored
  • Sunday

  • 24668 08/08/99 23:51 Ransome shift summary
  • 24667 08/08/99 22:01 x. jiang spectrometer window at run 878
  • 24645 08/08/99 18:26 x. jiang spectrometer screen t run 868
  • 24624 08/08/99 11:55 Calarco status report
  • 24609 08/08/99 07:55 M Jones optics replay at delta = 0%
  • 24606 08/08/99 07:08 M Jones Replay of optics 2% delta
  • 24605 08/08/99 06:45 M Jones optics at -4% delta
  • 24600 08/08/99 02:44 A. Deur eP energy measurement
  • 24599 08/08/99 01:25 M Jones Optics run plan
  • Saturday

  • 24598 08/07/99 23:56 Ransome Shift Summary
  • 24596 08/07/99 21:56 Saha Arc Energy Measurement
  • 24590 08/07/99 18:11 Saha Arc Integral measrement
  • 24588 08/07/99 17:38 Simon spectrometer screen at run 846
  • 24583 08/07/99 16:23 X. jiang 4:00 pm noticed that fpp 7.1.16 has been tripped
  • 24580 08/07/99 16:15 gilman real/random e,e'K coincidence rates
  • 24579 08/07/99 16:11 Calarco shift summary
  • 24575 08/07/99 15:05 R.Gilman current during e,e'K runs
  • 24574 08/07/99 15:01 Calarco spot
  • 24570 08/07/99 14:11 kuss sorry, the # is 838
  • 24568 08/07/99 14:07 kuss run #837
  • 24564 08/07/99 13:35 Bimbot Cryotarget
  • 24556 08/07/99 12:24 gilman/calar Moller quad supply
  • 24555 08/07/99 10:14 gilman hall a access
  • 24549 08/07/99 04:11 S. Strauch Shift summary
  • Friday

  • 24529 08/06/99 23:55 Markowitz shift summary
  • 24521 08/06/99 23:24 X. jiang 23:25 fpp chamber 4 hv 7.1.16 tripped again
  • 24514 08/06/99 23:07 X. jiang fpp ch4 hv 7.1.16 tripped again,
  • 24511 08/06/99 22:54 X. jiang put in 6% radiator
  • 24506 08/06/99 22:30 X. jiang FPP ch4 HV, channel 7.1.16 draws 3.6 uA
  • 24503 08/06/99 22:14 A. Deur eP energy measurement
  • 24494 08/06/99 21:27 X. jiang daq hung up at the end
  • 24492 08/06/99 21:13 X. jiang FPP threshold voltage and low voltage supply
  • 24491 08/06/99 21:10 X. jiang Get ready for 120 degree c.m. setting
  • 24486 08/06/99 20:28 kuss HRSH angle
  • 24485 08/06/99 20:09 kuss HRSH angle
  • 24482 08/06/99 17:37 X. Jiang FPP ch4 hv channel change.
  • 24478 08/06/99 16:53 R. Michaels DAQ tests runs 792 - 802
  • 24450 08/06/99 15:50 R. Michaels self-timing peaks
  • 24449 08/06/99 15:41 gilman ROC troubles run 788
  • 24436 08/06/99 11:51 Louis BIMBOT Cryotarget change
  • 24430 08/06/99 10:11 Saha Arc Energy measurement - Aug 5, 99
  • 24429 08/06/99 09:21 Saha Arc Energy Measurement - Aug 2, 99
  • 24428 08/06/99 09:15 gilman ch 4 hv run 780
  • 24425 08/06/99 08:25 R.Gilman IC levels
  • 24423 08/06/99 07:47 Meekins Shift Summary
  • 24419 08/06/99 07:17 meekins E arm dipole trip
  • 24396 08/06/99 02:50 Crovelli tripofflast
  • 24389 08/06/99 00:49 Crovelli 8116trip again
  • 24383 08/06/99 00:02 markowitz shift summary
  • Thursday

  • 24371 08/05/99 17:41 Markowitz data check
  • 24370 08/05/99 17:03 Markowitz fpp chamber 4
  • 24367 08/05/99 16:32 kuss dying klystron
  • 24366 08/05/99 16:17 X. Jiang/ R. day shift activities
  • 24364 08/05/99 16:15 markowitz energy measurements
  • 24360 08/05/99 15:49 jpchen moved 15 cm LD2 target in
  • 24351 08/05/99 14:54 gilman FPP efficiency
  • 24350 08/05/99 14:29 gilman plans for statistics on this d(g,p)n point
  • 24349 08/05/99 12:26 jpchen Online VDC effeciency calculation
  • 24348 08/05/99 12:22 jpchen Target moved to empty at 10 am.
  • 24346 08/05/99 09:33 gilman ch 4 trip no effect
  • 24342 08/05/99 08:08 M Jones gamma D runs owl shift
  • 24338 08/05/99 07:23 S. Strauch FPP HV 8.1.16 and events of interest
  • 24336 08/05/99 06:53 Adaq Fraction of useable events
  • 24333 08/05/99 06:27 M Jones Things to change in gammap_online when kinematics change
  • 24330 08/05/99 05:12 Meekins Target IOC reboot problem
  • 24326 08/05/99 03:45 Crovelli Doors reset
  • 24306 08/05/99 01:33 Crovelli t3plotradiator
  • 24298 08/05/99 00:21 Crovelli tripoutlast
  • 24294 08/05/99 00:02 Markowitz shift summary
  • Wednesday

  • 24289 08/04/99 23:39 markowitz test of radiator thickness vs t3 rates
  • 24274 08/04/99 22:14 markowitz spot for radiator test
  • 24272 08/04/99 20:36 R.Gilman radiator levels
  • 24271 08/04/99 20:26 R.Gilman Radiator test current 25microA
  • 24270 08/04/99 19:56 R.Gilman Radiator test empty target
  • 24268 08/04/99 19:27 kuss target moving
  • 24263 08/04/99 16:01 Arun Saha MCC time estimate
  • 24262 08/04/99 15:43 gilman fpp ch 4 trip
  • 24261 08/04/99 15:32 Arun Saha Controlled access
  • 24260 08/04/99 13:19 gilman helicity scalers
  • 24259 08/04/99 13:15 kuss manual for Ion Chamber StripTool
  • 24258 08/04/99 13:09 gilman ch 4 trip
  • 24257 08/04/99 12:24 gilman access into hadron arm
  • 24256 08/04/99 12:02 Arun Saha Conrolled access
  • 24255 08/04/99 11:55 X. Jiang VDC tracking and front FPP tracking compare
  • 24253 08/04/99 10:17 R. Michaels some DAQ checks
  • 24252 08/04/99 10:15 Arun Saha MCC request to tune injector
  • 24243 08/04/99 09:24 Arun Saha Bob's request for DAQ
  • 24241 08/04/99 09:11 R. Michaels pedestal suppression on ROC3
  • 24238 08/04/99 08:48 Arun Saha OTR carbon foil in.
  • 24236 08/04/99 08:37 Arun Saha Arc Integral measurement
  • 24234 08/04/99 08:30 M Jones fpp time offsets
  • 24233 08/04/99 08:26 gilman spot size
  • 24220 08/04/99 06:49 Ron Gilman tripwire again
  • 24217 08/04/99 06:40 Meekins Bremsstrahlung Radiator thickness tests
  • 24211 08/04/99 06:27 Ron Gilman wire trip 620
  • 24198 08/04/99 05:43 M JOnes Fpp Hv 8.1.16 tripped
  • 24197 08/04/99 05:33 Meekins hardron arm HV's
  • 24196 08/04/99 05:31 Meekins Electron Arm HV's
  • 24195 08/04/99 05:21 Eugene Moller results
  • Tuesday

  • 24189 08/03/99 21:42 markowitz spot check for moller running
  • 24179 08/03/99 18:35 M Jones auto-filtering on adaqs3
  • 24176 08/03/99 17:12 X. Jiang magnet screen
  • 24175 08/03/99 17:11 X. Jiang The 3" carbon door is in.
  • 24174 08/03/99 16:48 kuss HRSE aerogel settings
  • 24173 08/03/99 16:43 Ron Gilman access activities done
  • 24172 08/03/99 16:37 Ron Gilman. poolarized beam from injector
  • 24169 08/03/99 16:13 R. Michaels trigger download (and a reminder HOW)
  • 24168 08/03/99 16:00 Ron Gilman. e-arm Aerogel
  • 24166 08/03/99 15:57 Ron Gilman. beam status
  • 24165 08/03/99 15:52 R. Michaels trigger work done
  • 24159 08/03/99 15:23 kuss MCC status
  • 24158 08/03/99 11:35 A Deur eP energy measurement
  • 24157 08/03/99 11:29 R. Michaels scaler checks / new setup
  • 24156 08/03/99 11:22 Ron Gilman. RF time signal
  • 24148 08/03/99 06:13 Markowitz electron arm floor angle
  • 24147 08/03/99 06:05 Markowitz change settings
  • 24140 08/03/99 04:29 Nasseripour trigger timing
  • 24139 08/03/99 04:28 markowitz pion trigger timing
  • 24098 08/03/99 00:00 X. jiang run 657 beam heating test. 10 uA
  • Monday

  • 24097 08/02/99 23:52 kuss start of boiling studies
  • 24091 08/02/99 23:38 x. jiang run 654, spot check
  • 24090 08/02/99 23:36 x. jiang arc energy measurement
  • 24089 08/02/99 23:34 kuss boiling parameters
  • 24081 08/02/99 21:54 X. jiang magnet setting during run 651.
  • 24080 08/02/99 21:52 X. Jiang run 651, c12 target, raster off,
  • 24077 08/02/99 21:24 X. jiang run 650, LH2 target 15 cm cell, sieve slit
  • 24074 08/02/99 21:00 X. jiang magnet screen during run 649.
  • 24070 08/02/99 20:26 X. jiang FPP HV 8.1.16 tripped.
  • 24067 08/02/99 20:03 X. jiang run 646 ,4cm dummy cell in,
  • 24059 08/02/99 19:22 X. jiang run 643 dummy 15 cm target in
  • 24056 08/02/99 19:01 X. jiang Magnet screen during run 642.
  • 24053 08/02/99 18:35 X. jiang Magnet screen during run 641
  • 24039 08/02/99 13:26 X. Jiang eset FPP HV 8.1.16 to -1875.0
  • 24038 08/02/99 13:01 jpchen RCS complete, access to get ready for optics
  • 24036 08/02/99 12:58 X. Jiang Checked FPP high Voltage at 1:00 pm,
  • 24035 08/02/99 08:43 Adaq test
  • 24034 08/02/99 08:32 R. Michaels download correct trigger (RCS)
  • Sunday

  • 24031 08/01/99 16:03 J. Hines photon spectrometer
  • 24029 08/01/99 13:33 L. Taub angle change
  • 24028 08/01/99 13:19 R. Michaels raster current was unplugged !
  • 24024 08/01/99 11:22 kuss UPS software crash reported
  • 24021 08/01/99 10:11 kuss run #630
  • 24019 08/01/99 10:06 kuss LH2 15cm cell
  • 24018 08/01/99 09:21 kuss run #629
  • 24015 08/01/99 09:05 kuss LH2 4cm cell with radiator
  • 24013 08/01/99 09:03 kuss loop 3 fan speed
  • 24001 08/01/99 02:37 P. Markowitz hadron arm angle